ELECTRON SPIN RESONONANCE (ESR) STUDIES OF NEON MATRIX-ISOLATED (HF)_{2}^{−} AND (DF)_{2}^{−} DIMER RADICAL ANIONS: VALENCE OR DIPOLE-BOUND?

\underline{BEVAN ELLIOTT} and LON B. KNIGHT, JR., Department of Chemistry, Furman University, Greenville, SC 29613.

Electron spin resonance (ESR) spectroscopic studies of the hydrogen halide dimer anion radicals (HF)_{2}^{−} and (DF)_{2}^{−} are presented. These paramagnetic species were created from the parent HF and DF gases using a rare gas microwave discharge, and trapped in neon matrices at 4 K. Whether the matrix-isolated radicals are conventional valence anions or have dipole-bound character is not fully known, but (HF)_{2}^{−} exists in the gas phase only as a dipole bound anion. This study of the dimer anions represents the first observation of (HF)_{2}^{−} outside the gas phase or isolated in any matrix, and the first observation of (DF)_{2}^{−} by any means.